OpenStudy (anonymous):

Please Help! Will medal! xxx Marcus is eating a meal that is made up of an entrée that is 45 calories per ounce and a side dish that is 30 calories per ounce. He wants his meal to total 510 calories. Let e represent the number of ounces of the entrée that he eats, and let s represent the number of ounces of the side dish that he eats. Which function relates the number of ounces of the side dish that he eats to the number of ounces of the entrée that he eats? A. s= 2/3e - 17 B. s= 3/2e - 11 1/3 C. s= -2/3e + 11 1/3 D. s= -3/2e + 17

OpenStudy (anonymous):

I think it's A or C.

OpenStudy (anonymous):

Okay Thanks @Musics_life c:

OpenStudy (lyrae):

510 = 45e + 30s Solve for s.

OpenStudy (anonymous):

I got ..c? ( im sorry im super bad at math.) ;-;@Lyrae

OpenStudy (lyrae):

510 = 45e + 30s 30s = -45e + 30 \[s = - \frac{ 45 }{ 30 }e + \frac{ 510 }{ 30 }\] s = -3/2 e + 17

OpenStudy (lyrae):

Typo 30s = -45e + 30 Should be 30s = -45e + 510

OpenStudy (anonymous):

Ohhhh!! okay I get it now! cx thanks!

OpenStudy (lyrae):

Great! Yw :)